OpenStudy (andrewkaiser333):

Simplify 2^5-3^4/2^7-3^2

OpenStudy (igreen):

\(\dfrac{2^5-3^4}{2^7-3^2}\) Subtract the exponents. 5- 7 = ? 4 - 2= ?

OpenStudy (andrewkaiser333):

yes that is what the equation looks like but need help with answer

OpenStudy (igreen):

5 - 7 = ? 4 - 2 = ?

OpenStudy (andrewkaiser333):

this one is stumping me lol i cant figure it out

OpenStudy (andrewkaiser333):

-2 2

OpenStudy (andrewkaiser333):

but then what is next that is as far as i got

OpenStudy (igreen):

Okay, that gives us: \(\dfrac{2^{-2}}{3^2}\) Okay, is that it or do you have to simplify it more? @AndrewKaiser333

OpenStudy (andrewkaiser333):

more

OpenStudy (igreen):

\(2^{-2} = 0.25\) \(3^2 = 9\) \(\dfrac{0.25}{9}\) Can you divide that?

OpenStudy (igreen):

Oh my bad that gives us: \(2^{-2} - 3^2\) \(0.25 - 9\)

OpenStudy (igreen):

Can you subtract that? @AndrewKaiser333

OpenStudy (isaiah.feynman):

@iGreen the division rules for exponents cannot work for a problem like this.

OpenStudy (andrewkaiser333):

can i just get the answer please i am having to hurry the kid is oging to wake up and i am gonna have to feed him

OpenStudy (isaiah.feynman):